Suppose supply and demand curves are:

0 0
Answer #1

Setting QD=QS

28-2P=-4+2P

28+4=2P+2P

32=4P

Equilibrium P=32/4 = 8

Equilibrium Q=28-2(8) = 12

Maximum reservation price = 28/2 = 14

Minimum reservation price = 4/2 = 2

When a tax of $2 is imposed on the sellers,the supply curve will shift upwards by $2 at each price level

Price consumers pay = 9

Price producers receive = 7

New equilibrium quantity = 10

Deadweight loss = 0.5*[(12-10)*(9-7)] = 2